The height of SHS is normally distributed with a mean of = 150 cm and a standard deviation of = 10 cm.


a) Sketch a normal curve that describes this distribution.


b) Approximately what percent of these students have a height greater than 170 cm?




1
Expert's answer
2022-03-19T02:40:04-0400

a)



b)

"\mu=150,\sigma=10"

"P(X>170)=1-P(X<170)"

"=1-P(Z<\frac{170-150}{10})=1-P(Z<2)"

"=1-0.9772=0.0228"

"=2.28" %